HETAV PATEL

Grade 11,

Find the equivalent fraction of 36 / 48 with numerator 9

Find the equivalent fraction of 36 / 48 with
 numerator 9 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given numerator = 9
∴9 / D = 36 / 48
D × 36 = 9 × 48
D = (9 × 48) / 36
D = 12
Hence, the equivalent fraction is 9 / 12

Thanks